A factory has decided to diversify its activities. The data collected for the sales and production departments are summarized below: Potential demand exists for two products A and B. Market can absorb any quantity of A, whereas the share of B for this organization is expected to be not more than 400 units per month. For every three units of B produced there is one unit of a by - product which sells at K3 per unit and only 100 units of this by - product can be sold per month. Contribution per unit of products A and B is expected to be K6 and K8 respectively. These products require three different processes and the time required per unit of product is given in the table below: Process Product Product B Available hours 900 600 1200 III Find the product mix to optimize the contribution by using Simplex Method.


1
Expert's answer
2022-01-11T12:45:20-0500

x1 is units of product A

x2 is units of product B


maximize contribution:

"Z=6x_1+8x_2"

subject to:

"2x_1+3x_2\le 900" - process 1

"x_1+2x_2\le 600" - process 2

"2x_1+2x_2\le 1200" - process 3

"x_2\le 400"


After introducing slack variables:

Max Z=6x1+8x2+0S1+0S2+0S3+0S4

subject to

2x1+3x2+S1=900

x1+2x2+S2=600

2x1+2x2+S3=1200

x2+S4=400

and 

x1,x2,S1,S2,S3,S4≥0



Negative minimum Zj-Cj is -8 and its column index is 2. So, the entering variable is x2.

Minimum ratio is 300 and its row index is 2. So, the leaving basis variable is S2.

∴ The pivot element is 2.

Entering =x2, Departing =S2, Key Element =2





Negative minimum Zj-Cj is -2 and its column index is 1. So, the entering variable is x1.

Minimum ratio is 0 and its row index is 1. So, the leaving basis variable is S1.

∴ The pivot element is 0.5.

Entering =x1, Departing =S1, Key Element =0.5





Negative minimum Zj-Cj is -2 and its column index is 4. So, the entering variable is S2.

Minimum ratio is 150 and its row index is 2. So, the leaving basis variable is x2.

∴ The pivot element is 2.

Entering =S2, Departing =x2, Key Element =2




Since all Zj-Cj≥0

Hence, optimal solution is arrived with value of variables as :

x1=450,x2=0

Max Z=2700